A principal at a specific school thinks that his students are above average intelligence. He collects a random sample of IQ scores from 25 students from his school. His sample has a mean of 112. The mean population IQ is 100 with a standard deviation of 15. Does he have enough evidence to support his claim?
Answer #1

Null Hypothesis H0: u= 100

Alternative Hypothesis Ha: 001 +11

Alpha= 0.05

Z= 112-100/15/sqrt(25)

Z= 12/15/5

Z= 12/3

Z= 4

Z critical= 1.65

Since Z cal is greater than Z critical therefore significant.

Decision: Reject H0

Conclusion: We have sufficient evidence to show that the students are above average intelligence.